A projectile with mass 201.0 kg is launched straight up from Earth's surface with an initial speed 10.7 km/s. What is the maxium height of the projectile, as measured from the center of the Earth? Answer in units of Re (g= 10 m/s^2, Re= 6.371 x 10^6 m) The answer is 4.9 but how, Please and Thankyou
Answer #1


maximum height from the csurface of earth = u^2/(2*g) = 10700^2/(2*10)= 5724500 m

om the center of earth = Re+5724500 = (6.371*10^6)+(5.724500*10^6)= 12.0955 m..
in units of Re = 12.0955/(6.371*10^6) = 1.89 *Re